# Balanced Data Sampling for Language Model Training with Clustering

### v1

@@ Abstract
Data plays a fundamental role in the training of Large Language Models (LLMs). While attention has been paid to the collection and composition of datasets, determining the data sampling strategy in training remains an open question. Most LLMs are trained with a simple strategy, random sampling. However, this sampling strategy ignores the unbalanced nature of training data distribution, which can be sub-optimal. In this paper, we propose ClusterClip Sampling to balance the text distribution of training data for better model training. Specifically, ClusterClip Sampling utilizes data clustering to reflect the data distribution of the training set and balances the common samples and rare samples during training based on the cluster results. A repetition clip operation is introduced to mitigate the overfitting issue led by samples from certain clusters. Extensive experiments validate the effectiveness of ClusterClip Sampling, which outperforms random sampling and other cluster-based sampling variants under various training datasets and large language models.
@@ 2 Related Works
The composition of pre-training data is a determinant of LLM performance. Efforts are made to collect and utilize domain mixtures of pre-training data to improve LLM performance Longpre et al. (2023); Shen et al. (2023); Nijkamp et al. (2023). Brown et al. (2020); Chung et al. (2022); Du et al. (2022); Azerbayev et al. (2023); Touvron et al. (2023a, b) exploit manually designed domain composition weights by small-scale pre-training experiments. However, we focus on cluster-based data sampling approaches, which do not rely on manual selection of data composition. In addition, existing works also explore the algorithms for searching optimal domain weights. Several methods involve training auxiliary models, either proxy models or reference models to determine the composition weights Xie et al. (2023a); Fan et al. (2023); Mindermann et al. (2022). Furthermore, some works concentrate on sample-level data selection by introducing model-based metrics to measure sample weights, which include importance score Xie et al. (2023b), perplexity Xia et al. (2023), and gradients Marion et al. (2023). Different from these works, the cluster-based sampling strategies utilize off-the-shelf embedding models and semantic-based cluster-level data sampling.
@@ 4.1 Baselines
Inspired by recent practice of training domain-specific language models Rozière et al. (2023); Azerbayev et al. (2023), we want the model to learn general abilities before acquiring specific domain knowledge and skills. Thus, G2S initiates by uniform sampling from each cluster. If a particular cluster is exhausted, sampling from that cluster ceases until the entire dataset has been traversed. By doing so, the model learns diverse and general samples before concentrating on some specific clusters.
@@ 4.2 Training Datasets
We use the base-sized Jina Embeddings 2111https://huggingface.co/jinaai/jina-embeddings-v2-base-en Günther et al. (2023) with mean pooling to generate text embeddings. For Open-Orca, we do not differentiate between data generated by GPT-4 and GPT-3.5 and mix them for clustering and training, which reflects the results in a more realistic scenario with mixed instruction data quality. We concatenate the inputs and outputs and truncate the result text into 1024 tokens for embedding computation. Then, we run K-Means with cosine distance on generated embeddings over 300 iterations to obtain 2000 clusters. For Proof-Pile-2, we combined and shuffled different sub-domains of the mixture together and sampled 10B tokens as our training set. This approach allows us to explore the model behavior when changing data sampling strategies in the context of continual training on data mixtures with multiple domains. After obtain the document embeddings, we set the number of iterations and clusters of K-Means to 300 and 100 to obtain the clusters, respectively.
@@ 5.2 Analysis ClusterClip Sampling in Training
To find out how the ClusterClip Sampling affect the model performance as the training goes, we massively evaluate the intermediate checkpoints of models trained with different sampling strategies. The supervised fine-tuning results are shown in Figure 4 and the pre-training results are shown in Figure 5. We also present the data distribution of the training set in Figure 6 to see the connection of the data distribution and the results of these sampling methods.
@@
It is worth-noting that the training order of the samples may need to be carefully scheduled depending on different datasets, based on the results of G2S and S2G Sampling. It can be related to domain-specific learning or curriculum learning Bengio et al. (2009); Hacohen and Weinshall (2019); Rozière et al. (2023), which can be investigated in future work. And the cluster-based sampling methods are worth further exploration in this area.
@@ 5.3 Ablation of ClusterClip Sampling
We aim to investigate the impact of the number of clusters in the application of ClusterClip Sampling methods. We conduct training of Llama2-7B on Proof-Pile-2 datasets with different numbers of clusters, following other setups in the main experiments. The results are shown in Figure 9, in which the ClusterClip sampling consistently outperforms Uniform sampling by a large margin on the MMLU benchmarks, even though the number of clusters is changing from 100 to 1000, respectively. Moreover, the number of clusters does not affect the performance much under the same cluster-based sampling methods. Thus, the ClusterClip sampling is not sensitive to the number of clusters.
Figure 9: MMLU Accuracy of different sampling methods when changing the number of clusters on Proof-Pile-2 dataset.
@@ 6 Conclusion
We leveraged data clustering to find semantic clusters of training data and utilize the cluster information for data sampling. There is a risk that the cluster-based sampling aggregates the bias or toxic content in the dataset. However, in this work, we utilize open-sourced datasets to train the LLMs, including Open-Orca Lian et al. (2023) and Proof-Pile-2 Azerbayev et al. (2023). These datasets are specifically collected and filtered to avoid toxic and safety issues. When applying the cluster-based sampling strategies for new datasets, data cleaning and filtering methods can be used to remove the harmful content before clustering and training.

### v2

@@ Abstract
Data plays a fundamental role in the training of Large Language Models (LLMs). While attention has been paid to the collection and composition of datasets, determining the data sampling strategy in training remains an open question. Most LLMs are trained with a simple strategy, random sampling. However, this sampling strategy ignores the unbalanced nature of training data distribution, which can be sub-optimal. In this paper, we propose ClusterClip Sampling to balance the text distribution of training data for better model training. Specifically, ClusterClip Sampling utilizes data clustering to reflect the data distribution of the training set and balances the common samples and rare samples during training based on the cluster results. A repetition clip operation is introduced to mitigate the overfitting issue led by samples from certain clusters. Extensive experiments validate the effectiveness of ClusterClip Sampling, which outperforms random sampling and other cluster-based sampling variants under various training datasets and large language models  [^1] Our code is released at https://github.com/choosewhatulike/cluster-clip.
@@ 2 Related Works
The composition of pre-training data is a determinant of LLM performance. Efforts are made to collect and utilize domain mixtures of pre-training data to improve LLM performance Longpre et al. (2023); Shen et al. (2023); Nijkamp et al. (2023). Brown et al. (2020); Chung et al. (2022); Du et al. (2022); Azerbayev et al. (2023); Touvron et al. (2023a, b) exploit manually designed domain composition weights by small-scale pre-training experiments. However, we focus on cluster-based data sampling approaches, which do not rely on manual selection of data composition. In addition, existing works also explore the algorithms for searching optimal domain weights. Several methods involve training auxiliary models, either proxy models or reference models to determine the composition weights Mindermann et al. (2022); Xie et al. (2023a); Fan et al. (2023); Suzuki et al. (2023). Furthermore, some works concentrate on sample-level data selection by introducing model-based metrics to measure sample weights, which include importance score Xie et al. (2023b), perplexity Xia et al. (2023), and gradients Marion et al. (2023). Different from these works, the cluster-based sampling strategies utilize off-the-shelf embedding models and semantic-based cluster-level data sampling.
@@ 4.1 Baselines
Inspired by recent practice of training domain-specific language models Rozière et al. (2023); Azerbayev et al. (2023), we want the model to learn general abilities before acquiring specific domain knowledge and skills. Thus, G2S is initiated by uniform sampling from each cluster. If a particular cluster is exhausted, sampling from that cluster ceases until the entire dataset has been traversed. By doing so, the model learns diverse and general samples before concentrating on some specific clusters.
@@ 4.2 Training Datasets
We use the base-sized Jina Embeddings $2$https://huggingface.co/jinaai/jina-embeddings-v2-base-en Günther et al. (2023) with mean pooling to generate text embeddings. For Open-Orca, we do not differentiate between data generated by GPT-4 and GPT-3.5 and mix them for clustering and training, which reflects the results in a more realistic scenario with mixed instruction data quality. We concatenate the inputs and outputs and truncate the result text into 1024 tokens for embedding computation. Then, we run K-Means with cosine distance on generated embeddings over 300 iterations to obtain 2000 clusters. For Proof-Pile-2, we combined and shuffled different sub-domains of the mixture together and sampled 10B tokens as our training set. This approach allows us to explore the model behavior when changing data sampling strategies in the context of continual training on data mixtures with multiple domains. After obtaining the document embeddings, we set the number of iterations and clusters of K-Means to 300 and 100 to obtain the clusters, respectively.
@@ 4.3 Training Setups
4.4 Cost Comparison
We calculate the costs of clustering and training to figure out whether the clustering operation is a good investment when training a language model at scale. As shown in Table 1, the cost of clustering is relatively low compared with the large language model (LLM) training, because the model to generate document embedding is very small (100M in our experiments), and the K-means algorithm is very fast by paralleling on GPUs. The Proof-Pile-2 is much faster because it contains much longer documents, e.g. Arxiv papers and web pages, and the clustering only considers the first 1024 tokens. Thus, the clustering cost mainly depends on the number of samples in datasets, instead of the actual tokens.
Moreover, clustering is a one-time investment, and we may train many models of varying scales once the training data is clustered, which can further amortize the clustering costs. To further speed up the clustering and lower the memory usage, one can use more efficient methods and shrink the embedding size as mentioned in  Yamada et al. (2021); Kusupati et al. (2022).
Dataset Tokens Samples Avg. Length Embedding Storage
Open-Orca 1.8B 4.2M 429 tokens 14GB
Proof-Pile-2 13.4B 2.7M 4963 tokens 9GB
Dataset Embedding Cost Kmeans Cost Total Clustering Cost Training 7B Cost
Open-Orca 6 GPU hours 0.3 GPU hour 6.3 GPU hours 384 GPU hours
Proof-Pile-2 5.5 GPU hours 0.3 GPU hour 5.8 GPU hours 1216 GPU hours
Table 1: Clustering and Training Costs of Datasets.
@@ 5.2 Analysis ClusterClip Sampling in Training
To find out how the ClusterClip Sampling affects the model performance as the training goes on, we massively evaluate the intermediate checkpoints of models trained with different sampling strategies. The supervised fine-tuning results are shown in Figure 4 and the pre-training results are shown in Figure 5. We also present the data distribution of the training set in Figure 6 to see the connection between the data distribution and the results of these sampling methods.
@@
It is worth noting that the training order of the samples may need to be carefully scheduled depending on different datasets, based on the results of G2S and S2G Sampling. It can be related to domain-specific learning or curriculum learning Bengio et al. (2009); Hacohen and Weinshall (2019); Rozière et al. (2023), which can be investigated in future work. And the cluster-based sampling methods are worth further exploration in this area.
@@ 5.3 Ablation of ClusterClip Sampling
We aim to investigate the impact of the number of clusters in the application of ClusterClip Sampling methods. We conduct training of Llama2-7B on Proof-Pile-2 datasets with different numbers of clusters, following other setups in the main experiments. The results are shown in Figure 9, in which the ClusterClip sampling consistently outperforms Uniform sampling by a large margin on the MMLU benchmarks, even though the number of clusters is changing from 50 to 1000, respectively. Moreover, the number of clusters does not affect the performance much under the same cluster-based sampling methods. Thus, the ClusterClip sampling is not sensitive to the number of clusters.
Figure 9: MMLU Accuracy of different sampling methods when changing the number of clusters on the Proof-Pile-2 dataset.
@@ 6 Conclusion
We leveraged data clustering to find semantic clusters of training data and utilize the cluster information for data sampling. There is a risk that the cluster-based sampling aggregates the bias or toxic content in the dataset. However, in this work, we utilize open-sourced datasets to train the LLMs, including Open-Orca Lian et al. (2023) and Proof-Pile-2 Azerbayev et al. (2023). These datasets are specifically collected and filtered to avoid toxic and safety issues. When applying the cluster-based sampling strategies for new datasets, data cleaning, and filtering methods can be used to remove the harmful content before clustering and training.
Acknowledgements
This work was supported by the National Key Research and Development Program of China (No.2022ZD0160102). The computations in this research were performed using the CFFF platform of Fudan University.